OpenStudy (anonymous):

Graph the Solutions of each system; 1.{y>x-5 {3x+y|x-3|+1

OpenStudy (shamil98):

Solve the systems. y > x - 5 3x + y <= -2 substitute for y 3x + x - 5 <= -2 4x -5 <= -2 add 5 to both sides 4x <= 3 x <= 3/4 plug in for y y > 3/4 -5 y > -17/4

OpenStudy (anonymous):

What is the LCD? 1/45ab^2 and 5a/18b

OpenStudy (shamil98):

(3/4, -17/4) are your solution coordinates.

OpenStudy (shamil98):

Try doing number two on your own, by following my explanation.

OpenStudy (anonymous):

Okay Im going to try and do that I will let you know what I get! and thanks so much for your help :)